OpenStudy (anonymous):

What is the slope-intercept form of the equation y-7=-(5/2) (x+4)? y-2=-5/2 (x+2) y+7=-x+(5/2) y=-(4/7)+2 y=(-(5/2)x) -3

OpenStudy (anonymous):

Slope-intercept form is \(y = mx + b\) So you just have to make it look like that. \[y - 7 = -\frac{5}{2}(x + 4)\]\[y = -\frac{5}{2}x - \frac{5}{2}\times4+7\]\[y = -\frac{5}{2}x - 3\]

OpenStudy (texaschic101):

y - 7 = -5/2(x + 4) -- distribute through the parenthesis y - 7 = -5/2x - 20/2 y - 7 = -5/2x - 10 -- now add 7 to both sides y = -5/2x - 10 + 7 -- combine like terms y = -5/2x - 3